OpenStudy (firejay5):

Solve each system of equation by elimination: Show work and steps. 31. 3x + 5y = 6 2x - 4y = -7

OpenStudy (jackellyn):

Multiply the first one by 2 and the second by 3

OpenStudy (firejay5):

do you mean -3 @jackellyn

OpenStudy (hana88):

no the first 1 should be multiplied by -2

OpenStudy (hana88):

or what fire ay said

OpenStudy (firejay5):

why is it * by -2

OpenStudy (hana88):

because you need to make the first term the opposite of the second term to cancel it out

OpenStudy (firejay5):

did I do that right @jackellyn

OpenStudy (jackellyn):

that is right @hana88 and @Firejay5 one does have to be negative

OpenStudy (anonymous):

(3x+5y=6) multiply by -2 you get -6x-10y=-12 (2x-4y=-7) multiply by 3 you get 6x-12y=-21 by solving the above 2 equations you get y= -33/-22=3/2 substituting y = 3/2 in the 2nd eqn. you get x=-1/2